OpenStudy (anonymous):

Rationalize 5/sqrt(7) I forgot what it means to rationalize lol.

OpenStudy (anonymous):

\[5\div \sqrt{7}\]

OpenStudy (anonymous):

means make the denominator a rational number by writing \[\frac{5}{\sqrt{7}}\times \frac{\sqrt{7}}{\sqrt{7}}=\frac{5\sqrt{7}}{7}\]

OpenStudy (anonymous):

/facepalm Can't believe I forgo that. Thanks!